OpenStudy (anonymous):

how do you solve this equation -12+2v= -3(8v+4). Can anyone please show me the steps to solve these kinds of equations?

OpenStudy (anonymous):

Collect together all your v's on one side of the equals and all your numbers on the other so -12+2v= -3(8v+4) -12 + 2v = -24v -12 2v +24v = -12 +12 = 0 26v = 0 so v= 0